OpenStudy (anonymous):

Simplify the expression. Assume that all variables represent positive numbers. Write your answer without using negative exponents. (125u^6v^6)^1/3

OpenStudy (anonymous):

first write each of them in terms cubic powers. 125 = 5^3 u^6=(u^2)^3 v^6=(v^2)^3 so you can write 125u^6v^6 as (5u^2v^2)^3 so finally (125u^6v^6)^1/3 = ((5u^2v^2)^3)^1/3 = 5u^2v^2

OpenStudy (anonymous):

UR AWESOME!!! THANKS!!